# Optimal probe states for single-mode quantum target detection in arbitrary object reflectivity

### v1

@@ Abstract
Quantum target detection (QTD) utilizes nonclassical resources to enable radar-like detection for identifying reflecting objects in challenging environments, surpassing classical methods. To fully exploit the quantum advantage in QTD, determining the optimal probe states (OPSs) across various detection parameters and gaining a deeper understanding of their characteristics are crucial. In this study, we identified the single-mode continuous-variable OPSs for arbitrary object reflectivity using optimization algorithms. Our findings suggest that OPSs are non-Gaussian states in most reflectivity scenarios, with exceptions under specific conditions. Furthermore, we provide a comprehensive physical interpretation of the observed phenomena. This study offers a tool for identifying OPSs along with a clear physical interpretation. It also contributes to further advancements towards optimal multi-mode QTD, which holds the potential for broad applications in quantum sensing and metrology.
@@ I Introduction
Quantum target detection (QTD) is a critical task within the field of quantum sensing. In the context of QTD, a reflecting target is situated amidst a lossy and noisy environment, like typical radar detection scenarios. The probe emits a quantum signal to the target area, receives the reflected signal, and further analyzes its quantum nature to ascertain the presence of the target Pirandola et al. (2018); Giovannetti et al. (2011); Degen et al. (2017); Polino et al. (2020). To enhance the effectiveness of QTD, Lloyd initially introduced discrete variable (DV) entanglement into QTD, known as quantum illumination (QI) Lloyd (2008); Shapiro (2020). Subsequently, the concept of QI was applied to various types of quantum states as probes, including Gaussian states in continuous-variable (CV) systems Guha and Erkmen (2009); Pirandola and Lloyd (2008); Shapiro and Lloyd (2009); Tan et al. (2008); Nair and Gu (2020), and even non-Gaussian states Zhang et al. (2014); Sanz et al. (2017). Furthermore, a series of experimental demonstrations of QTD based on QI was presented in the optical frequency region Zhang et al. (2013, 2015) and even in the microwave region Barzanjeh et al. (2020); Assouly et al. (2023), laying the foundation for the implementation of type-III quantum lidar/radar Lanzagorta (2012); Burdge et al. (2009).
@@
In the aforementioned work, these studies are all based on a known probe quantum state as the foundation for research. The discussion revolves around the quantum advantage of this quantum state under stringent probing conditions (high loss and noisy environment). For example, in Shapiro and Lloyd (2009), a two-mode squeezed vacuum (TMSV) is discussed as the probe quantum state and achieves a quantum advantage of 6 dB. However, an intriguing question arises: are these discussed quantum states the optimal probe states (OPSs)? Is achieving a higher advantage using other types of quantum states under the same probing conditions possible? In DV systems, the issue of OPSs is addressed in Yung et al. (2020), where the maximally entangled state is identified as the OPS. Subsequent experiments demonstrate that using the maximally entangled state as the probe state can approach the theoretical limit of the Helstrom bound Xu et al. (2021). In CV systems, much research focuses on verifying the OPSs under conditions of low reflection targets, with results showing that the TMSV is the OPS for such probing conditions De Palma and Borregaard (2018); Bradshaw et al. (2021). Additionally, under the condition of without entanglement (i.e., single mode) in the probing source, the near-coherent state is theoretically discussed as the OPS Bradshaw et al. (2021); Spedalieri and Pirandola (2021).
@@
Although studies on OPSs have provided solutions under strict conditions (such as extremely high losses or low target reflectivity), there has been no clear discussion on CV OPSs for targets with finite or even high reflectivity. These non-extreme conditions more accurately reflect real-world detection scenarios. Moreover, the study of the arbitrary reflectivity of objects could also be analogous to various detection tasks. For example, when dealing with targets exhibiting high reflectivity, the situation resembles scenarios where the optical field is reflected back and detected under conditions of low loss. This implies weak interaction between the target and the probing light, a condition commonly encountered in measurements involving biological samples Taylor and Bowen (2016); Allen et al. (2020); Whittaker et al. (2017).
@@
In order to understand OPSs under global reflectivity, in this study, we employ optimization algorithms to determine the single-mode OPS that minimizes the error probability of QTD under arbitrary detection conditions. Drawing from the observed behavior of OPS under different conditions, we also delve into OPS from various physical perspectives to offer reasonable interpretations of its behavior. Our results show that in most cases, single-mode OPSs are non-Gaussian, but under specific background noise and target reflectivity conditions, OPS can revert to coherent states. Our work addresses OPS under arbitrary target reflectivity conditions, providing a comprehensive solution for QTD. The developed approach can be further extended to explore higher mode quantum states for optimal probe states with quantum entanglement.
@@
The paper is structured as follows: In Sec.II, we present the theoretical model of quantum target detection, covering aspects such as error probability estimation, beam splitter (BS) model, and optimal probe states. Subsequently, in Sec.III, we provide numerical results and discuss the underlying physical concepts. To present our findings systematically, we initially examine an athermal environment (noise-free environment) in Sec.III.A. Building upon the insights gained in this section, we then extend our discussion to a more general scenario of a thermal environment in Sec.III.B. In Sec.IV, we summarize the results of this study. Finally, we conclude our work in Sec.V.
@@ II Theoretical model
To estimate the error probability $P_{err}$, the states $\hat{\rho}_{0}$ and $\hat{\rho}_{1}$ must be calculated. Here, we utilize a BS model to analogy the lossy and noisy detection scenario, as shown in Fig.1 (b), and further calculate the states $\hat{\rho}_{0}$ and $\hat{\rho}_{1}$. In the BS model, a quantum transmitter inputs a quantum state of $\hat{\rho}_{pr}$ as a probe to a BS as the target with a reflectivity $r$. The noisy environment is simulated by a mixed thermal state $\hat{\rho}_{env}$ coupled into another mode of BS (see Fig.1). The thermal state is given by
@@
Based on the model above, one can understand that the error probability $P_{err}$ is a function of the variable set of $\{c_{n}\}$ with the constraints of Eq.6 and Eq.7. Once a specific set of $\{c_{n}\}$ which lets the $P_{err}$ become minima, the set of $\{c_{n}\}$, the corresponding probe states, is called optimal probe states. To find the OPSs, we regard the error probability $P_{err}$ as a target function to perform the Sequential Quadratic Programming (SQP) Boggs and Tolle (1995). After processing the algorithm, the output set of $\{c_{n}\}$ or the OPSs are obtained.
@@ III.1 Athermal environment ($\bar{n}_{env}=0$)
where $P_{m}$ is the probability that measured $m$ photon, $\rho_{pr,nn}$ is the photon number distribution of the initial probe state, and $r$ is the reflectivity or efficiency of photon-number-counting measurement. For $m=0$, it presents the probability of $\hat{\rho}_{1}$ occupying the vacuum state, which is contributed by the decayed nonzero photon states. Thus, we can obtain that from Eq.10 as
@@
where $p_{n}=\bar{n}-\left\lceil\bar{n}\right\rceil+1$ and $\left\lceil x\right\rceil=\min\{n\in\mathbb{Z}|x\leq n\}$.
Figure 3: Comparison of methods of photon statistics (PS, Eq.11) and density matrix (DM, Eq.1). (a) $\bar{n}=1$; (b) $\bar{n}=1.25$; (c) $\bar{n}=1.5$; (d) $\bar{n}=1.75$; (e) $\bar{n}=2$. The reflectivity is set at $r=0.99$ for all tests. The fidelity between the distributions calculated by the two methods is consistently higher than 0.999.
@@
is calculated and shown in Fig.2 (b). The corresponding Wigner function and density matrix of OPS in the low-reflectivity case ($r=10^{-3}$) are also plotted in Fig.2 (c) and (d), respectively. It is clear to see the OPSs in the $r\rightarrow 0$ region tend to be in a coherent state since the fidelity approach to 1. The above results are consistent with and have been demonstrated in Bradshaw et al. (2021) using the method of Lagrange multipliers.
@@
Figure 4: (a) Comparison of methods of phase overlapping (PO, Eq.15) and density matrix (DM, Eq.1) with different mean photon numbers. The fidelity between the distributions calculated by the two methods is consistently higher than 0.999. (b) Demonstration of Wigner functions of vacuum state ($\rho_{0}$, dash line) and received state ($\rho_{1}$, solid line and color map) in the case of target present.
@@
Figure 5: (a) Quantum advantage of OPSs in the case of $\bar{n}_{env}=\{0.04;0.1;0.2\}$ which are represented by the color of yellow, gray, and black, respectively, as indicated by the label in the figure. The mean photon number is set at 0.04. (b) The standard deviation (SD) ratio of photon number (square) and phase (circle) between OPSs and coherent state is presented. (c) The coherence of OPSs with a series of $\bar{n}_{env}$ as shown by the label in figure.
@@ III.2 Noisy environment ($\bar{n}_{env}\neq 0$)
In Fig.5 (a), the QA approaches 0 in the $r\rightarrow 0$ region. Moreover, in Fig.5 (b), the ratio of the variance of the photon number distribution and the phase distribution both tends towards 1, indicating that OPSs behave approach coherent states in this scenario. This outcome mirrors the discussion in an athermal environment. In a noisy setting, the phase distribution of thermal states remains constant. Therefore, the optimal discrimination method is assigning a defined phase to $\hat{\rho}_{recv}$. In this case, Eq. 15 can be applied, resulting in OPSs behaving as coherent states even in the presence of noise in the $r\rightarrow 0$ region.
@@
Another crucial discovery is that when $0<r<r_{T}$, the SD ratio indicates that OPSs have a more precise phase distribution than coherent states (and higher variance in the photon number distribution), implying that OPSs are phase-squeezed states within this region. To comprehend this characterization, one should realize that this situation is quite different from the case of an athermal environment. For $n_{env}>\bar{n}$, the $\hat{\rho}_{env}$ now contributes unavoidable noise photons on the diagonal terms of the state $\hat{\rho}_{recv}$ not only in the region of $r\rightarrow 0$ but also at a finite reflectivity to a certain degree. Thus, the OPSs cannot always follow the strategy used in the athermal case (PNSSs as probe states). To optimally distinguish between $\hat{\rho}_{env}$ and $\hat{\rho}_{recv}$, the off-diagonal terms (the coherence) of the probe states now play a crucial role in providing the difference between $\hat{\rho}_{env}$ and $\hat{\rho}_{recv}$, further resulting in the phase-squeezed state in the $0<r<r_{T}$ region.
To validate the aforementioned idea, we introduce the size of off-diagonal elements to quantify the coherence, denoted as $\mathcal{C}$, of the OPSs and further analyze its relation with the reflectivities. The coherence of the quantum state $\hat{\rho}$ is defined as Baumgratz et al. (2014),
@@
In Fig.5, with the increase in $r$, we now elucidate and consolidate the behaviors and properties of OPSs across the entire range of $r$. In the region $0<r<r_{T}$, as $r$ increases, the contribution of the photon number distribution from the thermal states gradually loses influence on $\hat{\rho}_{recv}$. Thus, the optimal strategy for distinguishing between $\hat{\rho}_{recv}$ and $\hat{\rho}_{env}$ starts to shift from mainly using coherence to a combination of both photon number distribution and coherence. This results in the SD ratio of OPSs gradually tending towards 1. Finally, when $r=r_{T}$, the OPSs revert to coherent states. When $r>r_{T}$, $\hat{\rho}_{recv}$ start to be dominated by $\hat{\rho}_{pr}$. Therefore, the behavior of OPSs is true to PNSSs, as demonstrated in Fig.5 (b).

### v2

@@ Abstract
Quantum target detection (QTD) utilizes nonclassical resources to enable radar-like detection for identifying reflecting objects in lossy and noisy environments, surpassing the detection performance achieved by classical methods. To fully exploit the quantum advantage in QTD, determining the optimal probe states (OPSs) across various detection parameters and gaining a deeper understanding of their characteristics are crucial. In this study, we employ optimization algorithms to identify the single-mode continuous-variable OPSs for entire range of target reflectivity. Our findings suggest that OPSs are non-Gaussian states in most reflectivity scenarios, with exceptions under specific conditions. Furthermore, we provide a comprehensive physical interpretation of the observed phenomena. This study offers a tool for identifying OPSs along with a clear physical interpretation. It also contributes to further advancements towards optimal multi-mode QTD, which holds the potential for broad applications in quantum sensing and metrology.
@@ I Introduction
Quantum target detection (QTD) is a critical task within the field of quantum sensing. In the context of QTD, a reflecting target is situated amidst a lossy and noisy environment, like typical radar detection scenarios. The probe emits a quantum signal to the target area, receives the reflected signal, and further analyzes its quantum nature to ascertain the presence of the target Pirandola et al. (2018); Giovannetti et al. (2011); Polino et al. (2020). To enhance the effectiveness of QTD, Lloyd initially introduced discrete variable (DV) entanglement into QTD, known as quantum illumination (QI) Lloyd (2008); Shapiro (2020). Subsequently, the concept of QI was applied to various types of quantum states as probes, including Gaussian states in continuous-variable (CV) systems Guha and Erkmen (2009); Pirandola and Lloyd (2008); Shapiro and Lloyd (2009); Tan et al. (2008); Nair and Gu (2020), and even non-Gaussian states Zhang et al. (2014); Sanz et al. (2017). In 2013, the first experimental realization of quantum illumination was presented Lopaeva et al. (2013) based on the protocol proposed by Lloyd and S. Tan et al. Lloyd (2008); Tan et al. (2008). Furthermore, a series of experimental demonstrations of QTD based on QI was presented in the optical frequency region Zhang et al. (2013, 2015) and even in the microwave region Barzanjeh et al. (2020); Assouly et al. (2023); Chang et al. (2019); Luong et al. (2019); Bourassa and Wilson (2020); Livreri et al. (2022, 2023); Hosseiny et al. (2022), laying the foundation for the implementation of type-III quantum lidar/radar Lanzagorta (2012); Burdge et al. (2009); Torromé and Barzanjeh (2023); Karsa et al. (2023); Torromé et al. (2020); Al-Moahmmed (2020). In more practical application scenarios, quantum radar systems must advance beyond previous quantum illumination protocols, which merely identify the presence of a target within a specific space. Recently, there has been further discussion on the three-dimensional localization of targets by quantum radar. Positioning within three-dimensional space by utilizing the protocol of Gaussian beam entangled photons in the frequency domain was first proposed in Maccone et al. (2023), which further demonstrates a $\sqrt{N}$ times quantum enhancement over the unentangled case when $N$ photons are used. Moreover, achieving more precise azimuth resolution of targets through the use of dual-receiver quantum radar protocols has also been discussed Wu et al. (2023).
@@
In the aforementioned work, these studies are all based on a known probe quantum state as the foundation for research. The discussion revolves around the quantum advantage of this quantum state under stringent probing conditions (high loss and noisy environment). For example, in Shapiro and Lloyd (2009), a two-mode squeezed vacuum (TMSV) is discussed as the probe quantum state and achieves a quantum advantage of 6 dB. However, an intriguing question arises: are these discussed quantum states the optimal probe states (OPSs)? Is achieving a higher advantage using other types of quantum states under the same probing conditions possible?
In DV systems, the issue of OPSs is addressed in Yung et al. (2020), where the maximally entangled state is identified as the OPS. Subsequent experiments demonstrate that using the maximally entangled state as the probe state can approach the theoretical limit of the Helstrom bound Xu et al. (2021). In CV systems, much research focuses on verifying the OPSs under conditions of low reflection targets. In Bradshaw et al. (2021), Mark et al. employ mathematical methods (Lagrange multiplier) to investigate the OPSs for quantum target detection under extreme conditions (high noise and low target reflectivity or high-loss channels) for both single-mode (without entanglement) and two-mode (involving entanglement) detection scenarios. They conclude that under single-mode and extreme detection conditions, the OPS approximates a coherent state, while for two-mode scenarios, the OPS corresponds to TMSV, which is consistent with the results presented in De Palma and Borregaard (2018). Furthermore, in the case of given probe states, optimizing the parameters of the probe state’s properties is also an essential topic for QTD. In Spedalieri and Pirandola (2021), the study focuses on using displaced squeezed states as probe states and investigates the joint optimal squeezing and displacement parameters of these states under various target reflectivities. It further demonstrates that these states can outperform coherent states with the same mean number of input photons.
@@
Although studies on OPSs have provided solutions under strict conditions (such as extremely high losses or low target reflectivity), there has been no clear discussion on CV OPSs for targets with finite or even high reflectivity. These non-extreme conditions more accurately reflect real-world detection scenarios. Moreover, the study of the arbitrary reflectivity of objects could also be analogous to various detection tasks. For example, when dealing with targets exhibiting high reflectivity, the situation resembles scenarios where the optical field is reflected back and detected under conditions of low loss. This implies weak interaction between the target and the probing light, a condition commonly encountered in measurements involving biological samples Taylor and Bowen (2016); Allen et al. (2020); Whittaker et al. (2017); Degen et al. (2017); Petrini et al. (2020).
@@
In this study, to comprehend OPSs under global reflectivity, we utilize optimization algorithms to identify the single-mode OPS that minimizes the error probability of QTD under arbitrary detection conditions. We demonstrate that the OPSs we identified exhibit higher identification performance than coherent state probes in QTD across arbitrary detection conditions. Drawing from the observed behavior of OPS under different conditions, we also delve into OPS from various physical perspectives to offer reasonable interpretations of its behavior. Our analysis shows that in most cases, single-mode OPSs are non-Gaussian, but under specific background noise and target reflectivity conditions, OPS can revert to coherent states. Furthermore, an important finding is that the probe-obtained information for distinguishing target absence or presence can be determined by the probe state’s phase or photon number distribution. This insight allows us to distinguish OPSs into two regimes based on target reflectivity: those dominated by phase-squeezed states and those dominated by photon number squeezed states (PNSSs), confirming the types of OPSs in given detection conditions. Our work provides a comprehensive understanding of OPSs under arbitrary target reflectivity conditions, offering insights for optimizing QTD techniques. The developed approach can be extended to explore higher mode quantum states for optimal probe states with quantum entanglement.
@@
The paper is structured as follows: In Sec.II, we present the theoretical model of quantum target detection, covering aspects such as error probability estimation, beam splitter (BS) model, and optimal probe states. Subsequently, in Sec.III, we provide numerical results and discuss the underlying physical concepts. To present our findings systematically, we initially examine an athermal environment (noise-free environment) in Sec.III.A. Building upon the insights gained in this section, we then extend our discussion to a more general scenario of a thermal environment in Sec.III.B. In Sec.IV, we summarize the results of this study. The optimal algorithm and the details for calculation are described in the Appendix.
@@ II Theoretical model
To estimate the error probability $P_{err}$, the states $\hat{\rho}_{0}$ and $\hat{\rho}_{1}$ must be calculated. Here, we utilize a BS model to analogy the lossy and noisy detection scenario, as shown in Fig.1 (b), and further calculate the states $\hat{\rho}_{0}$ and $\hat{\rho}_{1}$. In the BS model, a quantum transmitter inputs a quantum state of $\hat{\rho}_{pr}$ as a probe to a BS as the target with a reflectivity $r$. The noisy environment is simulated by coupling into another mode of BS (see Fig.1). To facilitate comparison with the majority of research on quantum target detection Guha and Erkmen (2009); Tan et al. (2008); Nair and Gu (2020); Zhang et al. (2014); Bradshaw et al. (2021), in this case, the environmental noise quantum state is considered a mixed thermal state $\hat{\rho}_{env}$. The thermal state is given by
@@
Based on the model above, one can understand that the error probability $P_{err}$ is a function of the variable set of $\{c_{n}\}$ with the constraints of Eq.6 and Eq.7. Once a specific set of $\{c_{n}\}$ which lets the $P_{err}$ become minima, the set of $\{c_{n}\}$, the corresponding probe states, is called optimal probe states. To find the OPSs, we regard the error probability $P_{err}$ as a target function to perform the Sequential Quadratic Programming (SQP) Boggs and Tolle (1995). After processing the algorithm (see Appendix), the output set of $\{c_{n}\}$ or the OPSs are obtained.
@@ III.1 Athermal environment ($\bar{n}_{env}=0$)
where $P_{m}$ is the probability that measured mm photon, $\rho_{pr,nn}$ is the photon number distribution of the initial probe state, and rr is the reflectivity or efficiency of photon-number-counting measurement. For $m=0$, it presents the probability of $\hat{\rho}_{1}$ occupying the vacuum state, which is contributed by the decayed nonzero photon states. Thus, we can obtain that from Eq.10 as
@@
where $p_{n}=\bar{n}-\left\lceil\bar{n}\right\rceil+1$, $\left\lceil x\right\rceil=\min\{n\in\mathbb{Z}|x\geq n\}$ and $\left\lfloor x\right\rfloor=\min\{n\in\mathbb{Z}|x\leq n\}$.
Figure 3: Comparison of the diagonal term of OPSs obtained by the methods of photon statistics (PS, Eq.11) and density matrix (DM, Eq.1).The mean photon number of probe state is set at (a) $\bar{n}=1$; (b) $\bar{n}=1.25$; (c) $\bar{n}=1.5$; (d) $\bar{n}=1.75$; (e) $\bar{n}=2$. The reflectivity is set at $r=0.99$ for all tests. The fidelity between the distributions calculated by the two methods is consistently higher than 0.999.
Figure 4: (a) Comparison of phase distribution of OPSs obtained by the methods of phase overlapping (PO, Eq.15) and density matrix (DM, Eq.1) in different mean photon numbers. The fidelity between the distributions calculated by the two methods is consistently higher than 0.999. (b) Demonstration of Wigner functions of vacuum state ($\rho_{0}$, dash line) and received state ($\rho_{1}$, color map) by using OPSs. The mean photon number of probe state is set at 1. The reflectivity is set at 0.01 for this figure.
@@
is calculated and shown in Fig.2 (b). The corresponding Wigner function and density matrix of OPS in the low-reflectivity case ($r=10^{-3}$) are also plotted in Fig.2 (c) and (d), respectively. It is clear to see the OPSs in the $r\rightarrow 0$ region tend to be in a coherent state since the fidelity approaches 1. The above results are consistent with and have been demonstrated in Bradshaw et al. (2021) using the method of Lagrange multipliers.
@@
Figure 5: (a) Quantum advantage of OPSs in the case of $\bar{n}_{env}=\{0.04;0.1;0.2\}$ which are represented by the color of yellow, gray, and black, respectively. The mean photon number is set at 0.04. (b) The standard deviation (SD) ratio of photon number (square) and phase (circle) between OPSs and coherent state is presented. (c) The coherence ratio between OPSs and coherent state with a series of $\bar{n}_{env}$ (as shown by the label in figure).
@@ III.2 Noisy environment ($\bar{n}_{env}\neq 0$)
In Fig.5 (a), the QA approaches 0 in the $r\rightarrow 0$ region. Moreover, in Fig.5 (b), the ratio of the variance of the photon number distribution and the phase distribution both tends towards 1, indicating that OPSs behave approach coherent states in this scenario. This outcome mirrors the discussion in an athermal environment. In a noisy setting, the phase distribution of thermal states remains constant. Hence, the optimal discrimination method involves assigning a defined phase to $\hat{\rho}_{recv}$, while Eq. 15 can also be applied in this case. This results in OPSs behaving as coherent states even in the presence of noise in the $r\rightarrow 0$ region.
@@
Another crucial discovery is that when $0<r<r_{T}$, the SD ratio indicates that OPSs have a more precise phase distribution than coherent states (and higher variance in the photon number distribution), implying that OPSs are phase-squeezed states within this region. To comprehend this characterization, one should realize that this situation is quite different from the case of an athermal environment. For $\bar{n}_{env}>\bar{n}$, the $\hat{\rho}_{env}$ now contributes unavoidable noise photons on the diagonal terms of the state $\hat{\rho}_{recv}$ not only in the region of $r\rightarrow 0$ but also at a finite reflectivity to a certain degree. Thus, the OPSs cannot always follow the strategy used in the athermal case (PNSSs as probe states). To optimally distinguish between $\hat{\rho}_{env}$ and $\hat{\rho}_{recv}$, the off-diagonal terms (the coherence) of the probe states now play a crucial role in providing the difference between $\hat{\rho}_{env}$ and $\hat{\rho}_{recv}$, further resulting in the phase-squeezed state in the $0<r<r_{T}$ region.
To validate the idea, we introduce the size of off-diagonal elements to quantify the coherence, denoted as $\mathcal{C}$, of the OPSs and further analyze its relation with the reflectivities. The coherence of the quantum state $\hat{\rho}$ is defined as Baumgratz et al. (2014),
@@
We now elucidate and consolidate the behaviors and properties of OPSs across the entire range of $r$. In Fig.5, in the region $0<r<r_{T}$, as $r$ increases, the contribution of the photon number distribution from the thermal states gradually loses influence on $\hat{\rho}_{recv}$. Thus, the optimal strategy for distinguishing between $\hat{\rho}_{recv}$ and $\hat{\rho}_{env}$ starts to shift from mainly using coherence to a combination of both photon number distribution and coherence. This results in the SD ratio of OPSs gradually tending towards 1. Finally, when $r=r_{T}$, the OPSs revert to coherent states. When $r>r_{T}$, $\hat{\rho}_{recv}$ start to be dominated by $\hat{\rho}_{pr}$. Therefore, the behavior of OPSs tends to PNSSs, as demonstrated in Fig. 5 (b).
Figure 6: Schematic diagram for the basic SQP algorithm. $\mathcal{L}$ is the Lagrangian for the problem, $\lambda_{k}$ and $\sigma_{k}$ are Lagrange multipliers. $\nabla^{2}_{xx}$ is Hessian matrix.